OpenStudy (anonymous):

lim 3x^7+x^2 x->-inf

OpenStudy (zarkon):

-infinity

OpenStudy (anonymous):

So you don't have to do the squeeze theorem or anything

OpenStudy (zarkon):

no

OpenStudy (anonymous):

Is this true for all polynomials or just this case

OpenStudy (zarkon):

with limits to infinity or -infinity?

OpenStudy (anonymous):

could explain both cases

OpenStudy (zarkon):

for a polynomial (with limits to - or + infinity) just look at what happens to the highest power term.

OpenStudy (zarkon):

\[\lim_{x\to-\infty}3x^7=-\infty\] therefore \[\lim_{x\to-\infty}3x^7+x^2=-\infty\]

OpenStudy (anonymous):

what if the power is negative 3x^-7+x^2 would 2 be the higher power or -7

OpenStudy (zarkon):

then it is not a polynomial and

OpenStudy (zarkon):

\[\lim_{x\to-\infty}3x^{-7}+x^2=\infty\]

OpenStudy (anonymous):

right polynomials don't negative powers. so what would you do in this case

OpenStudy (zarkon):

\[\lim_{x\to-\infty}3x^{-7}=0\] \[\lim_{x\to-\infty}x^2=\infty\]

myininaya (myininaya):

\[\lim_{x \rightarrow \infty}(\frac{3}{x^7}+x^2)=\lim_{x \rightarrow \infty}\frac{3+x^9}{x^7}=\lim_{x \rightarrow \infty}\frac{\frac{1}{x^7}}{\frac{1}{x^7}} \frac{3+x^9}{x^7}\] \[=\lim_{x \rightarrow \infty}\frac{\frac{3}{x^7}+\frac{x^9}{x^7}}{1}=\lim_{x \rightarrow \infty} \frac{0+\infty}{1}=\infty\]

myininaya (myininaya):

and of course i didn't mean to leave that lim x->infty there after i took the limit of each part

OpenStudy (anonymous):

I am sorry I am not getting it, thank you for your help

myininaya (myininaya):